OpenStudy (anonymous):

is it c? What is 2 to the power of 4 over 3 equal to? cube root of 8 square root of 8 cube root of 16 square root of 16

OpenStudy (anonymous):

@Agl202 no clue?

OpenStudy (johnweldon1993):

\[\large 2^{\frac{4}{3}} = \sqrt[3]{2^4}\] \[\large \sqrt[3]{2\times 2\times 2\times 2}\] \[\large \sqrt[3]{16}\] So C yes

OpenStudy (jdoe0001):

\(\large { a^{\frac{{\color{blue} n}}{{\color{red} m}}} \implies \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}}\qquad thus \\ \quad \\ 2^{\frac{{\color{blue} 4}}{{\color{red} 3}}}\implies ? }\)

OpenStudy (anonymous):

thanks @johnweldon1993